# Biomedical Image Segmentation: A Systematic Literature Review of Deep Learning Based Object Detection Methods

### v1

@@ 5.2 Articles related to One-stage Detection
The detailed descriptions of these models along with the types of imaging modalities, the types of diseases these models have been used for the detection, the performance evaluation metrices and quantitative measures are shown in Table 3.
@@ 5.2.1 Analysis of One Stage Detectors
Like two stage detection models, the modalities that have been targeted by one stage detection models include MRI, CT, PET, X-Rays, dermoscopic images, microscopic images, ultrasound images, endoscopic, colonoscopic, histopathology, fundus images, and few other modalities. Among the one stage detection models, YOLO and its different versions are the most powerful and frequently applied models. The most frequently YOLO version applied for medical image segmentation is YOLO version 3 followed by YOLO version 5, YOLO version 4, and YOLO version 8, respectively. Few authors have applied YOLO version 2 and YOLO version 7 as well whereas some authors have applied SSD, RetinaNet, and RefineNet for the segmentation of biomedical images but their frequency is very less. The quantitative measures of different performance evaluation parameters are shown in the last column of the table in which many variations can be observed. The variations in the results are due to different reasons including easy processing and implementation, less computational cost, high accuracy, simple architecture, and strength for handling complex datasets. All these factors lead YOLO and its versions to overperform other one stage models.
@@ 5.3 Articles Related to Point-Based Detection Models
There are some object detection models for image segmentation which have some different properties both from one-stage detection models and two-stage detection models. Keeping in consideration these properties, the models have been placed in a different category which has been named as advanced object detection models. The most common models included in this category include CornerNet, CornerNet-Lite, objects as points, and Foveabox. In literature, many authors have applied these object detection models for the segmentation of biomedical imaging. For example, Nawaz et al.  [170] applied CornerNet for the segmentation of brain MRI to analyse it for brain tumour detection. Similarly, Nawaz et al.  [171] applied a combined model of CornerNet and Fuzzy K-means for the segmentation of dermoscopic images to detect melanoma in the skin. Similarly, Wang et al.  [172] proposed the use of CornerNet for the segmentation of colonoscopic images to detect polyp. In another study, VASILEIOS ALEVIZOS  [173] proposed a backbone of CornerNet for the segmentation of brain MRI to detection brain tumour. ALADHAD and MAHUM  [174] proposed CornerNet of the extraction of informative features from the X-ray images to segment it for the detection of knee osteoarthritis.
@@
Sandhya et al  [175] presented CornerNet for the segmentation of MRI images to detect two types of malignant cancers namely low-grade glioma and high-grade glioma. Choi et al.  [176] suggested two deep learning models namely centernet and U-net for detection and segmentation of endoscopic images to detect gastrointestinal disease. Zheng et al  [177] proposed an improved version of Centernet for the segmentation of microscopic images to detect white blood cells in them. In the same manner, Jiang et al.  [178] proposed a modified version of CenterNet for the segmentation of colonoscopic images to detect polyp. In another work, Wang et al.  [179] suggested CenterNet for the segmentation of Optical Coherence Tomography Angiography (OCTA) images to detect retinal vascular bifurcation and crossover points. Similarly, Masood et al  [180] applied CenterNet for the segmentation of brain MRI images to detect tumour.
@@
Similarly, Albahli and Nazir  [181] suggested CenterNet for the segmentation of X-Ray images to classify and localize the diseases in them. In another study, Thangaval and Palanichamy  [182] applied CenterNet for the segmentation of CT images to detect pulmonary cancers. Jasitha and Pournami  [183] analysed both the faster R-CNN and CenterNet for the segmentation of histopathology images to detect Glomeruli. In another study, Nazir et al.  [184] applied CenterNet for the segmentation of retinal images to detect diabetic retinopathy. Shen et al.  [185] applied the combination of R-CNN, CenterNet, and deep snake for thyroid nodule detection in the ultrasound images. Mattikalli et al.  [186] developed an ensemble model of VFNet, FoveaBox, and RetinaNet for the detection of lesion in CT images. Similarly, Mathai et al.  [187] applied a hybrid model of VFNet, FoveaBox, FCOS, and DetectoRS for detection of lymph nodes in MRI. Sakunpaisanwari et al.  [188] different two stage detector models and one stage detector models including CornerNet and CenterNet for blood vessels detection in CT images. Durak et al.  [144] implemented different two stage detectors, one stage detectors, and other advanced detection models including CenterNet, for the detection of gastric polyps in the endoscopic images. Mathai et al.,  [189] applied an ensembled model of foveabox with other one stage detectors to detect lymphadenopathy in MRI images. Similarly, Celik et al.  [190] studied the role of different deep learning models including foveabox in the detection of lesions on panoramic radiographs. Cho et al.  [191] compared the performance of RetinaNet with cornernet and centernet in automatic tip detection of surgical instruments in biportal endoscopic spine surgery. In another work, Sakunpaisanwari et al.  [188] applied different deep learning models including centernet and cornernet for blood vessels detection in CT scans of lower extremities.
@@
Nguyen et al.  [192] applied different deep learning models including centernet and cornernet for the detection of glomeruli and nuclei in pathological images. In another work, Zheng et al.  [177] applied centernet for the detection of blood cells in the microscopic images. In another work, Su et al.  [193] presented a centernet based on heatmap pyramid structure for the detection of rib fracture in CT images of chest. Kitvimornat et al.  [194] applied faster R-CNN, RetinaNet, and Centernet for the detection and characterization of parasite eggs in the microscopic images. Similarly, Jiang et al.  [195] suggested the use of centernet along with Resnet 50 for feature extraction in the diagnosis of ultrasound images. In another attempt, Dai et al.  [196] suggested the use of YOLOv4 and Centernet for the diagnosis of breast ultrasound to detect lesion. Wang et al.  [172] applied few deep learning models including centernet for the detection of polyp segmentation in colonoscopy. Similarly, Ma et al.  [197] proposed a transformer based centernet for the liver tumor detection in CT images. The detailed descriptions of all these models along with other specifications are shown in Table 4.
@@
Table 4: Advanced detectors for biomedical imaging segmentation
@@ 5.3.1 Analysis of Point-based Detectors
Similarly, the advanced detectors have been applied for the segmentation of different biomedical imaging modalities including MRI, X-Rays, dermoscopic images, colonoscopic images, microscopic images, OCTA, CT images, histopathological images, fundus images, ultrasound images, endoscopic images, and radiographs. If their performance is deeply analysed, the highest accuracies have been achieved using both cornernet and centernet with similar results in few cases and few fluctuations with each other. In some cases, cornernet outperforms centerent whereas in other cases, centernet outperforms cornernet. The results obtained from Foveabox are not very satisfactory in terms of accuracy. The major reasons behind this phenomenon of different results include size of datasets used in the experimentations, running environment, models parameters’ tuning, data divisions into training and testing sets, data augmentation approach, data pre-processing and post processing mechanisms among the most important factors.
6 CHALLENGES AND FUTURE DIRECTIONS
There are many challenges associated with different types of biomedical images as well the models applied for their segmentitons. This section outlines all the challenges suggest future directions related to these challenges and future trends in biomedical image segmentation used for diseases detection.
@@ 6.1 Challenges with Biomedical Image Segmentation
The segmentation of biomedical image segmentation face many challenges due to the complex nature of images data, variability of images, high power computational requirements, insufficient data availability, and different hardware and software requirements among other factors. The first and foremost challenge for the segmentation of biomedical images is the process of accruacte annotation and labeling. This is due to the fact that segmentation is not an easy task like classification which requires few values in the output rather it requires images in the output for preparing training datasets. This is impossible without the involvement of field expert along with high quality software tools which needs additional time, efforts, and cost. Secondly, the complex structure of biomedical images due to various shapes, formats, organs anatomy, cells structures and tissues structures have very deep effect on the performance of segmentation models. Another challenging task associated with the biomedical image segmentation is the scale of biomedical images which ranges from smaller nuclei of a cell to lareger organs of human body e.g., hearts, lungs, kidney, and other organs. Developing robust segmentation models for this large scale differences is highly challenging.
Another important and considerable issue related to biomedical images is the class imbalance problem  [198]. This issue arises due to unavailability of sufficient instances for all the available classes of diseases being considered in the problem. This problem has a very bad impact on the performance of the segmentation models which leads to higher inaccuracy in the segmentation process and biases in the prediction. Similarly, all types of biomedical images are affected by different types of noises which adequately degrade the quality of images. Other factors affecting the quality of medical images include low contrast, distortions, low quality illumination, and body movements during .capturing images. All these types of unwanted activities have negative impact on biomedical images processing resulting in poor performance of the segmentation tasks. To address all of these challenges, we need effective, efficient, and powerful communication and collaborations among different experts including medical experts, computer experts, and scientists. This will reduce the barriers prevailing in the processing of biomedical images which will improve the process of segmentation.
@@ 6.3 Future Research Directions
4) Unsupervised and Semi-Supervised Approaches: Labeling and annotating biomedical images require significant resources, including time and effort. Unsupervised and semi-supervised learning approaches could alleviate some of these challenges. Models such as Self Organizing Maps (SOM), Generative Adversarial Networks (GANs), graph-based techniques, and autoencoders represent promising areas for future research.
@@
7) Hybrid approaches with large data: Exploring interdisciplinary approaches that combine biomedical imaging the data from other related domains such as bioinformatics, computational physics, and systems biology could also be explored to analyse the images based on different types of related data. For example, with the use of genetic data and integrating it in image segmentation could enhance the understanding of disease mechanisms at a molecular level, leading to more personalized and effective treatments.

### v2

@@ 5.2 Articles Related to One-Stage Detection
Similarly, different skin diseases namely melanoma, dysplastic nevus, and healthy skin have been identified with RetinaNet [68]. Deep RetinaNet has been used for the segmentation of dermoscopic images to identify the melanoma lesion [67]. A hybrid model of RetinaNet and ResNet was applied for the detection of skin cancer in dermoscopic images [66]. Furthermore, Bagheri et al.  [172] proposed three different object detection models for the identification of skin lesions in the dermoscopic images. The detailed descriptions of these models along with the types of imaging modalities, the types of diseases these models have been used for the detection, the performance evaluation metrices and quantitative measures are shown in Table 3.
@@ 5.2.1 Analysis of One Stage Detectors
Similar to two-stage detection models, the modalities targeted by one-stage detection models include MRI, CT, PET, X-rays, dermoscopic images, microscopic images, ultrasound, endoscopic, colonoscopic, histopathology, fundus images, and several other modalities. Among the one-stage detection models, YOLO and its various versions are the most powerful and frequently applied. The most commonly used YOLO version for medical image segmentation is YOLOv3, followed by YOLOv5, YOLOv4, and YOLOv8, respectively. Some authors have also applied YOLOv2 and YOLOv7, while others have used SSD, RetinaNet, and RefineNet for biomedical image segmentation, though their frequency of use is much lower. The quantitative measures of different performance evaluation parameters are shown in the last column of the table, where considerable variations can be observed. These variations result from several factors, including ease of processing and implementation, lower computational cost, high accuracy, simple architecture, and robustness in handling complex datasets. These factors contribute to YOLO and its versions outperforming other one-stage models.
@@ 5.3 Articles Related to Point-Based Detection Models
Some object detection models for image segmentation possess properties distinct from both one-stage and two-stage detection models. Due to these unique characteristics, they have been categorized separately as advanced object detection models or point-based detection models. The most common models in this category include CornerNet, CornerNet-Lite, Objects as Points, and FoveaBox. In the literature, many authors have applied these models to the segmentation of biomedical images. In this regard, CornerNet has been used for the segmentation of brain MRI to analyse it for brain tumour detection [173], for the segmentation of colonoscopic images to detect polyp [174], and for the extraction of informative features from the X-ray images to segment it for the detection of knee osteoarthritis [175]. It has also been used for the segmentation of MRI images to detect two types of malignant cancers namely low-grade glioma and high-grade glioma [176]. Another study proposed a backbone of CornerNet for the segmentation of brain MRI to detection brain tumour [177]. A combined model of CornerNet and Fuzzy K-means has been for the segmentation of dermoscopic images to detect melanoma in the skin [178].
@@
Recently, two deep learning models namely centernet and U-net have been used for detection and segmentation of endoscopic images to detect gastrointestinal disease  [179]. Another recent study proposed an improved version of Centernet for the segmentation of microscopic images to detect white blood cells in them [180]. Similarly, Jiang et al.  [181] proposed a modified version of CenterNet for the segmentation of colonoscopic images to detect polyp. In other studies the CenterNet has been used for the segmentation of Optical Coherence Tomography Angiography (OCTA) images to detect retinal vascular bifurcation and crossover points [182], and for the segmentation of brain MRI images to detect tumour [183].
@@
Similarly, Albahli and Nazir  [184] used CenterNet for the segmentation of X-Ray images to classify and localize the diseases in them. An other study applied CenterNet for the segmentation of CT images to detect pulmonary cancers [185]. Jasitha and Pournami  [186] analysed both the faster R-CNN and CenterNet for the segmentation of histopathology images to detect Glomeruli. Recently, Nazir et al.  [187] applied CenterNet for the segmentation of retinal images to detect diabetic retinopathy. In hybrid models, the combination of R-CNN, CenterNet, and deep snake for thyroid nodule detection in the ultrasound images [188]. Another hybrid model of VFNet, FoveaBox, FCOS, and DetectoRS for detection of lymph nodes in MRI [189]. Mattikalli et al.  [190] developed an ensemble model of VFNet, FoveaBox, and RetinaNet for the detection of lesion in CT images.
Sakunpaisanwari et al.  [191] different two stage detector models and one stage detector models including CornerNet and CenterNet for blood vessels detection in CT images. Another study implemented different two stage detectors, one stage detectors, and other advanced detection models including CenterNet, for the detection of gastric polyps in the endoscopic images [160]. Recently, an ensembled model of foveabox with other one stage detectors have been used to detect lymphadenopathy in MRI images [192]. Similarly, Celik et al.  [193] studied the role of different deep learning models including foveabox in the detection of lesions on panoramic radiographs. An other article [194] compared the performance of RetinaNet with cornernet and centernet in automatic tip detection of surgical instruments in biportal endoscopic spine surgery. Similarly, in an other study [191], different deep learning models including centernet and cornernet are used for blood vessels detection in CT scans of lower extremities.
@@
Nguyen et al.  [195] applied different deep learning models including centernet and cornernet for the detection of glomeruli and nuclei in pathological images. Furthermore, Zheng et al.  [180] applied centernet for the detection of blood cells in the microscopic images. In another work, a centernet based on heatmap pyramid structure for the detection of rib fracture in CT images of chest [196]. Faster R-CNN, RetinaNet, and Centernet have been used for the detection and characterization of parasite eggs in the microscopic images [197]. Similarly, Jiang et al.  [198] used centernet along with Resnet 50 for feature extraction in the diagnosis of ultrasound images. In another study use of YOLOv4 and Centernet for the diagnosis of breast ultrasound to detect lesion [199]. Recently, few deep learning models including centernet have been used for the detection of polyp segmentation in colonoscopy [174]. Transformer based centernet is also introduced which is used for the liver tumor detection in CT images [200]. The detailed descriptions of all these models along with other specifications are shown in Table 4.
@@
Table 4: Advanced or point-based detectors for biomedical imaging segmentation.
@@ 5.3.1 Analysis of Point-Based Detectors
Advanced (or Point-Based) detectors have been applied for the segmentation of various biomedical imaging modalities, including MRI, X-rays, dermoscopic images, colonoscopic images, microscopic images, OCTA, CT images, histopathological images, fundus images, ultrasound images, endoscopic images, and radiographs. When their performance is closely analyzed, the highest accuracies have been achieved using both CornerNet and CenterNet, with similar results in some cases and slight fluctuations between them. In certain instances, CornerNet outperforms CenterNet, while in others, CenterNet surpasses CornerNet. The results obtained from FoveaBox, however, are less satisfactory in terms of accuracy. The variations in these outcomes can be attributed to several factors, including the size of datasets used in the experiments, the running environment, model parameter tuning, data division into training and testing sets, data augmentation approaches, and data pre-processing and post-processing methods.
6 CHALLENGES & FUTURE DIRECTIONS
There are many challenges associated with different types of biomedical images as well the models applied for their segmentation. This section outlines all the challenges suggest future directions related to these challenges and future trends in biomedical image segmentation used for diseases detection.
@@ 6.1 Challenges with Biomedical Image Segmentation
Biomedical image segmentation faces many challenges including the complex nature of image data, variability in images, high computational requirements, insufficient data availability, and differing hardware, software requirements, an many other factors. The foremost challenge in this regard is the accurate annotation and labeling process. Unlike classification, which only requires a few output values, segmentation needs generating images as output for preparing training datasets. This task is not possible without the involvement of field experts and high-quality software tools, which require additional time, effort, and cost. Additionally, the complex structure of biomedical images—encompassing various shapes, formats, organ anatomies, cell structures, and tissue architectures—significantly impacts the performance of segmentation models. Another major challenge is the scale variation in biomedical images, ranging from the small nuclei of cells to large full organs like the heart, lungs, kidneys, and others. Developing robust segmentation models to accommodate these large-scale differences is particularly challenging.
Similarly, another important and considerable issue related to biomedical images is the class imbalance problem  [201]. This issue arises due to unavailability of sufficient instances for all the available classes of diseases being considered in the problem. This problem has a very bad impact on the performance of the segmentation models which leads to higher inaccuracy in the segmentation process and biases in the prediction. Similarly, all types of biomedical images are affected by different types of noises which adequately degrade the quality of images. Other factors affecting the quality of medical images include low contrast, distortions, low quality illumination, and body movements during .capturing images. All these types of unwanted activities have negative impact on biomedical images processing resulting in poor performance of the segmentation tasks. To address all of these challenges, we need effective, efficient, and powerful communication and collaborations among different experts including medical experts, computer experts, and scientists. This will reduce the barriers prevailing in the processing of biomedical images which will improve the process of segmentation.
@@ 6.3 Future Research Directions
4) Unsupervised and Semi-Supervised Approaches: Labeling and annotating biomedical images require significant resources, including time and effort. Unsupervised and semi-supervised learning approaches could alleviate some of these challenges. Models such as Self Organizing Maps (SOM), Generative Adversarial Networks (GANs), graph-based techniques, and auto-encoders represent promising areas for future research.
@@
7) Hybrid approaches with large data: Exploring interdisciplinary approaches that combine biomedical imaging the data from other related domains such as bio-informatics, computational physics, and systems biology could also be explored to analyse the images based on different types of related data. For example, with the use of genetic data and integrating it in image segmentation could enhance the understanding of disease mechanisms at a molecular level, leading to more personalized and effective treatments.
@@ 7 CONCLUSION
Acknowledgments
This work was supported by the Engineering and Physical Sciences Research Council (grant number: EP/X023826/1).